The reflection of the point (4,-13) in the S.L 5x+y+6=0 is the point??


Grade:

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear anand s

let the reflection is (h,k)

mid point of (h,k) and (4,-13) lies on straight line

and slope of line and the line joining by point(h,k) and (4,-13) are perpendicular

slope of the line =-5

slope of line  joining by point(h,k) and (4,-13)=(k+13)/(h-4)

so (k+13)/(h-4) * (-5) =-1

 5k +65 =h-4

or  h= 5k +69 ......1

 

and mid point lie on line

so 5(h+4)/2   + (k-13)/2   +6  = 0

 or 5(h+4)   + (k-13)   +12  = 0

     5h + k +19 =0..............2

solve above  two equation

point(-1 ,-14)


Please feel free to post as many doubts on our discussion forum as you can. If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin